### v1

@@ 6 Numerical Results
Next, we provide several proof-of-the-concept examples. [^1] Code available upon acceptance.
@@ 6.1 Joint Embedding of 3d Shapes
In the first example, we exploit EWλ to align and embed 3d shapes—the surfaces of objects in $\mathbb{R}^{3}$—into a joint space $(Z,d_{Z})$. For this, we interpret a 3d shape as mm-space $\mathbb{X}\coloneqq(X,d_{X},\xi)$, where $X$ is the surface, $d_{X}$ is the surface (or geodesic) distance, and $\xi$ is the uniform measure. Practically, $X$ is parametrized by the vertices of a triangular mesh, $d_{X}$ is approximated using Dijkstra’s algorithm Dijkstra (1959) on the corresponding graph, and $\xi$ is chosen as the discrete uniform measure. For the joint embedding of two given (discrete) shapes $\mathbb{X}_{i}$, $i=1,2$, we compute the 4-plan $\alpha$ in (14) using the discretization in Section 5 and Algorithm 1. Since the relaxation behind EWλ does not yield an isometry, but only a transport plan $\gamma_{i}=P_{X_{i}\times Z_{i},\sharp}\alpha$, we visualize the computed relaxed embeddings by the marginals $P_{Z_{i},\sharp}\gamma_{i}=P_{Z_{i},\sharp}\alpha$.
@@
Human Shapes
Since we are not limited to isometries, we next consider an experiment where isometric embeddings are not possible. More precisely, we embed human shapes from the FAUST dataset Bogo et al. (2014) into $\mathbb{R}^{2}$. For this, we choose $(Z,d_{Z})$ as an equispaced 60$\times$60 grid on $[0,1.3]^{2}\subset\mathbb{R}^{2}$ and apply Algorithm 1 with $\lambda=100$ and $\varepsilon=4\cdot 10^{-4}$. The obtained transport-based embeddings are shown in Figure 11. As comparison, we apply JMDS with $\lambda=10$ and $\varepsilon=4\cdot 10^{-4}$. The 2d representations clearly resemble the 3d human shapes, but JMDS splits some of the extremities.
@@ 6.2 Alignment of Feature Spaces
In the next example, we use our method to align different feature spaces occurring in real-world data. More precisely, we consider the genetyping—the determination of the corresponding cell class—of single cells from various measured modalities. Multimodal technologies as proposed in Cheow et al. (2016); Chen et al. (2019) are, however, uncommon such that there is a rising interest in embedding the corresponding features into a joint feature space Demetci et al. (2022); Liu et al. (2019); Chen et al. (2023). Following the experiments of Demetci et al. (2022) for $n$ single cells, our first aim is to assign the recorded features $X_{1}\coloneqq\{x_{1}^{1},\dots,x_{1}^{n}\}\subset\mathbb{R}^{d_{1}}$ from one modality to the features $X_{2}\coloneqq\{x_{2}^{1},\dots,x_{2}^{n}\}\subset\mathbb{R}^{d_{2}}$ of another modality, i.e., to recover the underlying one-to-one correspondence between $x_{1}^{j}$ and $x_{2}^{j}$. Our second goal consists in transferring a classifier from $X_{1}$ to $X_{2}$ in an unsupervised manner.
@@
1. We equip the recorded, uncorrelated features $X_{1}\subset\mathbb{R}^{d_{1}}$ and $X_{2}\subset\mathbb{R}^{d_{2}}$ with specific distances $d_{X_{i}}$ based on the path length in a nearest-ten-neighbor graph, see Cao et al. (2020); Demetci et al. (2022).
@@
For the first aim, the identification of the underlying correspondence, we identify $x_{1}^{j}$ with $x_{2}^{k}$ whose barycentric projection $\tilde{x}_{2}^{k}$ is closest to $\tilde{x}_{1}^{j}$. To quantify the quality of the pairing, we rely on the FOSCTTM (fraction of samples closer than the true match) score Liu et al. (2019). More precisely, FOSCTTM is defined by
@@
For the experiments, we employ the publicly available datasets [^2] https://rsinghlab.github.io/SCOT/data/ from Demetci et al. (2022), where SNAREseq consists of $d_{1}=19$ and $d_{2}=10$ features of $n=1047$ single cells, and scGEM of $d_{1}=34$ and $d_{2}=27$ features of $n=177$ specimens. The computed embedding of $X_{i}$ into $\mathbb{R}^{2}$ of our method and JMDS are visualized in Figure 14. For FOSCTTM and KKN-Acc, we additionally compare both methods with the alignment techniques UnionCom Cao et al. (2020) and SCOT Demetci et al. (2022) using the same distances on $X_{i}$. Note that UnionCom relies on GUMA and t-SNE, whereas SCOT relies on GW and MDS. The results are recorded in Table 1. The hyperparameters of all methods are chosen according to a grid search minimizing the FOSCTTM on a 10% validation split. A further experiment with different feature spaces for MNIST and FashionMNIST is given in Appendix C.
@@
Figure 14: Joint embedding of two feature spaces into $\mathbb{R}^{2}$ using our method and JMDS. (Top: first feature space, Bottom: second feature space) Both methods align the color-coded classes.
@@
${\operatorname{EW}}_{\lambda}$ (Ours) 0.000 0.877 0.000 0.679
JMDS Chen et al. (2023) 0.026 0.896 0.174 0.491
SCOT Demetci et al. (2022) 0.596 0.089 0.248 0.516
UnionCom Cao et al. (2020) 0.257 0.536 0.373 0.270
@@ 6.3 Alignment of Gaussian Mixture Models
In the final example, inspired by Salmona et al. (2024), we want to align Gaussian mixture models (GMMs). For this, we consider the space of 2d Gaussians $\mathcal{N}_{2}=\{\mathcal{N}(\mu,\Sigma)|\mu\in\mathbb{R}^{2},\Sigma\in\mathbb{R}^{2\times 2}\text{ spd.}\}$. } equipped with the Wasserstein distance. A GMM can now be interpreted as mm-space $\mathbb{X}=(\mathcal{N}_{2},\operatorname{W},\xi)$ with discrete measure $\xi\coloneqq\sum_{j=1}^{n_{1}}\xi_{j}\,\delta_{\mu_{j},\Sigma_{j}}$. For the experiment, we fit two GMMs to affine-transformed 2d datasets using the expectation-maximization algorithm. Here, we employ the datasets “Blobs” and “Moons” from scikit-learn Pedregosa et al. (2011), see Figure 17. For the alignment of the GMMs, we choose $Z\subset\mathcal{N}_{2}$, where the means form a 15$\times$15 grid over $[0,1]^{2}$ and we use a coarse grid over suitable matrices. Particularly, we consider matrices
@@
Figure 17: Aligenment of GMMs with respect to the Wasserstein distance.
@@ 7 Conclusion
We propose an unbalanced OT framework with GW marginal penalization which enables the joint embedding of two datasets based on pairwise intra-dataset distances. While our model can handle the aligned transfer into arbitrary metric spaces, it relies on their appropriate definition. In particular, due to computational restrictions, working on a grid in $\mathbb{R}^{d}$ is restricted to small dimensions $d$. As a future research direction, we are interested in developing a non-Euclidean free-support solver based on existing free-support GW barycenter algorithms Peyré et al. (2016); Vayer et al. (2020a); Beier & Beinert (2024). This may combine our EWλ approach with a non-Euclidean version of JMDS.
@@ Appendix A Proofs
The main idea of the proof follows the argumentation for the original embedded Wasserstein distance on Euclidean spaces in Salmona et al. (2024), which we adapt to arbitrary metric spaces. The definiteness follows directly from the definition of the isomorphic equivalence classes, the isometries in EW, and the definiteness of the Wasserstein distance. The symmetry is inherited from the Wasserstein distance too. It remains to show the triangle inequality. Without loss of generality, we take representatives $\mathbb{X}_{i}\coloneqq(Z,d_{Z},\xi_{i})$, $i=1,\dots,3$, such that $\xi_{i}\in\mathcal{P}(Z)$ possess full support. Since the inverse and composition of isometries are isometry, we have
@@
Proof of Proposition 3.2. Let $X\in\{X_{i}:i=1,2\}$. We equip the space of continuous functions $C(X,Z)$ with the topology of uniform convergence $d(f,g)\coloneqq\max_{x\in X}d_{Z}(f(x),g(x))$.
@@ 1. First, we we show that
is compact. Since $(Z,d_{Z})$ is compact, we see immediately that $\operatorname{Iso}(X,Z)$ is pointwise bounded. Moreover, since all functions in $\operatorname{Iso}(X,Z)$ are isometries the set is equicontinuous. By the Theorem of Arzelà–Ascoli Munkres (2000) (Thm. 47.1), this gives the relative compactness of $\operatorname{Iso}$. Further, $\operatorname{Iso}(X,Z)\subset C(X,Z)$ is closed and thus compact, since we have for any sequence of isometies $(I_{n})_{n}$ from $X$ to $Z$ converging to some $I\in C(X,Z)$ that
@@ 2. Next, we show that the functional
Thus we obtain the weak convergence $(I_{i,n})_{\sharp}\xi_{i}\rightharpoonup I_{i,\sharp}\xi_{i}$. Finally, due to the joint weak lower semi-continuity of the Wasserstein distance Ambrosio et al. (2005) (Lemma. 7.1.4), we get the desired lower semi-continuity of $F$.
@@
Proof of Proposition 3.3. Since $\mathcal{P}(Z\times Z)$ is weakly compact, the existence of a minimizer follows by Weierstrass’ theorem, see, e.g., Mordukhovich & Nam (2022) (Thm. 2.164), once we can show that the objective function
@@
and by continuity of the integrand and the fact that $(\gamma_{i,n}\otimes\gamma_{i,n})\rightharpoonup(\gamma_{i}\otimes\gamma_{i})$, see, e.g., Bogachev (2018) (Prop. 2.7.8) further
@@
Proof of Proposition 3.4. 1. First, we show that

### v2

@@ 6 Numerical Results
Next, we provide several proof-of-concept examples [^1] https://github.com/MoePien/RelaxedEmbeddedWasserstein. Additional parameter ablations are described in Supplement C.
@@ 6.1 Joint Embedding of 3d Shapes
In the first example, we exploit EWλ to align and embed 3d shapes—the surfaces of objects in $\mathbb{R}^{3}$—into a joint space $(Z,d_{Z})$. For this, we interpret a 3d shape as mm-space $\mathbb{X}\coloneqq(X,d_{X},\xi)$, where $X$ is the surface, $d_{X}$ is the surface (or geodesic) distance, and $\xi$ is the uniform measure. Practically, $X$ is parametrized by the vertices of a triangular mesh, $d_{X}$ is approximated using Dijkstra’s algorithm (Dijkstra, 1959) on the corresponding graph, and $\xi$ is chosen as the discrete uniform measure. For the joint embedding of two given (discrete) shapes $\mathbb{X}_{i}$, $i=1,2$, we compute the 4-plan $\alpha$ in (17) using the discretization in Section 5 and Algorithm 1. Since the relaxation behind EWλ does not yield an isometry, but only a transport plan $\gamma_{i}=P_{X_{i}\times Z_{i},\sharp}\alpha$, we visualize the computed relaxed embeddings by the marginals $P_{Z_{i},\sharp}\gamma_{i}=P_{Z_{i},\sharp}\alpha$.
@@
Human Shapes
Since we are not limited to isometries, we next consider an experiment where isometric embeddings are not possible. More precisely, we embed human shapes from the FAUST dataset (Bogo et al., 2014) into $\mathbb{R}^{2}$. For this, we choose $(Z,d_{Z})$ as an equispaced 60$\times$60 grid on $[0,1.3]^{2}\subset\mathbb{R}^{2}$ and apply Algorithm 1 with $\lambda=100$ and $\varepsilon=4\cdot 10^{-4}$. The obtained transport-based embeddings are shown in Figure 11. As comparison, we apply JMDS with $\lambda=10$ and $\varepsilon=4\cdot 10^{-4}$. The 2d representations clearly resemble the 3d human shapes, but JMDS splits some of the extremities. Since our method and JMDS can both be derived from the same functional, we additionally compare the achieved objectives, the Wasserstein distances between embeddings, and the GW distances between embeddings and input spaces. The experiment is repeated for each pairwise combination of the first ten shapes from the FAUST dataset and different $\lambda$. For comparison, the shapes are downsampled to 35 vertices, and both methods are regularized with $\varepsilon=0.001$. For our method, we set $Z$ according to a uniform 30$\times$30 grid on $[0,1.3]^{2}$. Results are shown in Table 1. In total, we here achieve better joint embeddings than JMDS both in terms of GW and Wasserstein.
GW2 W2
Ours ($\lambda=1)$ 0.0057$\pm$0.004 0.017$\pm$0.007
JMDS ($\lambda=1$) 0.0071$\pm$0.007 0.029$\pm$0.010
Ours ($\lambda=10$) 0.0048$\pm$0.004 0.027$\pm$0.016
JMDS ($\lambda=10$) 0.0057$\pm$0.005 0.046$\pm$0.023
Table 1: Mean and standard deviation of GW distance (between input shapes and embeddings) and of Wasserstein distance (between both embeddings) for joint 2d embeddings of the FAUST dataset via our method and JMDS.
@@ 6.2 Alignment of Feature Spaces
In the next example, we use our method to align different feature spaces occurring in real-world data. More precisely, we consider the genetyping—the determination of the corresponding cell class—of single cells from various measured modalities. Multimodal technologies as proposed in (Cheow et al., 2016; Chen et al., 2019) are, however, uncommon such that there is a rising interest in embedding the corresponding features into a joint feature space (Demetci et al., 2022; Liu et al., 2019; Chen et al., 2023). Following the experiments of (Demetci et al., 2022) for $n$ single cells, our first aim is to assign the recorded features $X_{1}\coloneqq\{x_{1}^{1},\dots,x_{1}^{n}\}\subset\mathbb{R}^{d_{1}}$ from one modality to the features $X_{2}\coloneqq\{x_{2}^{1},\dots,x_{2}^{n}\}\subset\mathbb{R}^{d_{2}}$ of another modality, i.e., to recover the underlying one-to-one correspondence between $x_{1}^{j}$ and $x_{2}^{j}$. Our second goal consists in transferring a classifier from $X_{1}$ to $X_{2}$ in an unsupervised manner.
@@
1. We equip the recorded, uncorrelated features $X_{1}\subset\mathbb{R}^{d_{1}}$ and $X_{2}\subset\mathbb{R}^{d_{2}}$ with specific distances $d_{X_{i}}$ estimated with the unsupervised SCOT routine (Demetci et al., 2022).
@@
For the first aim, the identification of the underlying correspondence, we identify $x_{1}^{j}$ with $x_{2}^{k}$ whose barycentric projection $\tilde{x}_{2}^{k}$ is closest to $\tilde{x}_{1}^{j}$. To quantify the quality of the pairing, we rely on the FOSCTTM (fraction of samples closer than the true match) score (Liu et al., 2019). More precisely, FOSCTTM is defined by
@@
Figure 14: Joint embedding of two feature spaces into $\mathbb{R}^{2}$ using our method and JMDS (Top: first feature space, Bottom: second feature space). Both methods align the color-coded classes.
@@
${\operatorname{EW}}_{\lambda}$ (Ours) 0.165 0.943 28.9s 0.208 0.610 1.7s
JMDS (Chen et al., 2023) 0.219 0.872 8.7s 0.230 0.616 0.3s
SCOT (Demetci et al., 2022) 0.402 0.447 76.7s 0.247 0.578 17.6s
UnionCom (Cao et al., 2020) 0.436 0.380 90.6s 0.392 0.465 1.5s
@@
For the experiments, we employ the publicly available datasets [^2] https://rsinghlab.github.io/SCOT/data/ from (Demetci et al., 2022), where SNAREseq consists of $d_{1}=19$ and $d_{2}=10$ features of $n=1047$ single cells, and scGEM of $d_{1}=34$ and $d_{2}=27$ features of $n=177$ specimens. The computed embedding of $X_{i}$ into $\mathbb{R}^{2}$ of our method and JMDS are visualized in Figure 14. For FOSCTTM and KKN-Acc, we additionally compare both methods with the alignment techniques UnionCom (Cao et al., 2020) and SCOT (Demetci et al., 2022) using the same distances on $X_{i}$. Note that UnionCom relies on GUMA and t-SNE, whereas SCOT relies on GW and MDS. The results are recorded in Table 2. The hyperparameters of all methods are chosen according to a grid search minimizing the FOSCTTM on a 10% validation split, see Supplement D. A further experiment with different feature spaces for MNIST and FashionMNIST is given in Supplement E.
@@ 6.3 Alignment of Gaussian Mixture Models
In the final example, inspired by (Salmona et al., 2024), we want to align Gaussian mixture models (GMMs). For this, we consider the space of 2d Gaussians $\mathcal{N}_{2}=\{\mathcal{N}(\mu,\Sigma)|\mu\in\mathbb{R}^{2},\Sigma\in\mathbb{R}^{2\times 2}\text{ spd.}\}$. } equipped with the Wasserstein distance, which yields a non-Euclidean geometry. A GMM can now be interpreted as mm-space $\mathbb{X}=(\mathcal{N}_{2},\operatorname{W},\xi)$ with discrete measure $\xi\coloneqq\sum_{j=1}^{n_{1}}\xi_{j}\,\delta_{\mu_{j},\Sigma_{j}}$. For the experiment, we fit two GMMs to affine-transformed 2d datasets using the expectation-maximization algorithm. Here, we employ the datasets “Blobs” and “Moons” from scikit-learn (Pedregosa et al., 2011), see Figure 17. For the alignment of the GMMs, we choose $Z\subset\mathcal{N}_{2}$, where the means form a 15$\times$15 grid over $[0,1]^{2}$ and we use a coarse grid over suitable matrices. Particularly, we consider matrices
@@
Figure 17: Alignment of GMMs with respect to the Wasserstein distance. The considered and computed GMMs are visualized via their density function on $\mathbb{R}^{2}$.
@@ 7 Conclusion
We propose an unbalanced OT framework with GW marginal penalization which enables the joint embedding of two datasets based on pairwise intra-dataset distances. While our model can handle the aligned transfer into arbitrary metric spaces, it relies on their appropriate definition. In particular, due to computational restrictions, working on a grid in $\mathbb{R}^{d}$ is restricted to small dimensions $d$. As a future research direction, we are interested in developing a non-Euclidean free-support solver based on existing free-support GW barycenter algorithms (Peyré et al., 2016; Vayer et al., 2020a; Beier & Beinert, 2025). This may combine our EWλ approach with a non-Euclidean version of JMDS.
From a theoretical point of view, the joint embedding using proposed variational formulation may be adapted to handle multiple input spaces. The main issue is hereby the selection of an appropriate generalization of the employed Wasserstein part. A possible choice would be to utilize only the Wasserstein distance between the embeddings of the $k$th and $(k+1)$st embedding. Since the employed multi-marginal Sinkhorn method linearly scales in the number of spaces, we expect that the resulting method remains numerically tractable for small numbers of spaces. The actual choice of the model and the numerical implementation of a multiple space variant, however, requires further studies.
Acknowledgements
M.P. acknowledges funding from the German Research Foundation (DFG, GRK2260 BIOQIC project 289347353).
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Proofs
The main idea of the proof follows the argumentation for the original embedded Wasserstein distance on Euclidean spaces in (Salmona et al., 2024), which we adapt to arbitrary metric spaces. The definiteness follows directly from the definition of the isomorphic equivalence classes, the isometries in EW, and the definiteness of the Wasserstein distance. The symmetry is inherited from the Wasserstein distance too. It remains to show the triangle inequality. Without loss of generality, we take representatives $\mathbb{X}_{i}\coloneqq(Z,d_{Z},\xi_{i})$, $i=1,\dots,3$, such that $\xi_{i}\in\mathcal{P}(Z)$ possess full support. Since the inverse and composition of isometries are isometry, we have
@@
Proof of Proposition 3.2
Let $X\in\{X_{i}:i=1,2\}$. We equip the space of continuous functions $C(X,Z)$ with the topology of uniform convergence $d(f,g)\coloneqq\max_{x\in X}d_{Z}(f(x),g(x))$.
@@ 1. First, we we show that
is compact. Since $(Z,d_{Z})$ is compact, we see immediately that $\operatorname{Iso}(X,Z)$ is pointwise bounded. Moreover, since all functions in $\operatorname{Iso}(X,Z)$ are isometries the set is equicontinuous. By the Theorem of Arzelà–Ascoli (Munkres, 2000) (Thm. 47.1), this gives the relative compactness of $\operatorname{Iso}$. Further, $\operatorname{Iso}(X,Z)\subset C(X,Z)$ is closed and thus compact, since we have for any sequence of isometies $(I_{n})_{n}$ from $X$ to $Z$ converging to some $I\in C(X,Z)$ that
@@ 2. Next, we show that the functional
Thus we obtain the weak convergence $(I_{i,n})_{\sharp}\xi_{i}\rightharpoonup I_{i,\sharp}\xi_{i}$. Finally, due to the joint weak lower semi-continuity of the Wasserstein distance (Ambrosio et al., 2005) (Lemma. 7.1.4), we get the desired lower semi-continuity of $F$.
@@
Proof of Proposition 3.3
Since $\mathcal{P}(Z\times Z)$ is weakly compact, the existence of a minimizer follows by Weierstrass’ theorem, see, e.g., (Mordukhovich & Nam, 2022) (Thm. 2.164), once we can show that the objective function
@@
and by continuity of the integrand and the fact that $(\gamma_{i,n}\otimes\gamma_{i,n})\rightharpoonup(\gamma_{i}\otimes\gamma_{i})$, see, e.g., (Bogachev, 2018) (Prop. 2.7.8) further
@@
Proof of Proposition 3.4
1. First, we show that
@@
Proof of Proposition 3.5
Without loss of generality, assume that $\pi_{n}$ converges to $\pi\in\Pi(\zeta_{1},\zeta_{2})$ with $\zeta_{i}\in\mathcal{P}(Z)$ weakly; otherwise take a convergent subsequence. Since $(Z,d_{Z})$ is compact, the diameter $M\coloneqq\sup_{z,z^{\prime}\in Z}d(z,z^{\prime})$ is finite. Using plans $\tilde{\pi}_{n}\in\Pi(\zeta_{1,n},\zeta_{2,n})$, where $\zeta_{i,n}\in\mathcal{P}(Z)$ are arbitrary GW approximations of $\mathbb{X}_{i}$, i.e., minimizers of (10), we estimate the relaxed embedded Wasserstein metric as follows:
$\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},P_{1,\sharp}\pi_{n}))+\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},P_{2,\sharp}\pi_{n}))\leq\frac{1}{\lambda_{n}}\operatorname{EW}_{\lambda_{n}}^{2}(\mathbb{X}_{1},\mathbb{X}_{2})$ (48)
$\leq\frac{1}{\lambda_{n}}\int_{Z\times Z}d_{Z}^{2}(z,z^{\prime})\,\mathrm{d}\tilde{\pi}_{n}(z,z^{\prime})+\sum_{i=1}^{2}\operatorname{GW}^{2}(\mathbb{X}_{i},(Z,d_{Z},P_{i,\sharp}\tilde{\pi}_{n}))\leq\frac{M^{2}}{\lambda_{n}}+\operatorname{GWA}(\mathbb{X}_{1})+\operatorname{GWA}(\mathbb{X}_{2}).$ (49)
Taking the limit $n\to\infty$, and exploiting (Beier & Beinert, 2025, Lem. I.1), we notice that the left-hand side becomes $\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},\zeta_{1}))+\operatorname{GW}^{2}(\mathbb{X}_{2},(Z,d_{Z},\zeta_{2}))$. Hence $\zeta_{i}$, $i=1,2$, have to be GW approximations of $\mathbb{X}_{i}$ on $(Z,d_{Z})$. $\Box$
Proof of Proposition 3.6
Without loss of generality, assume that $\pi_{n}$ converges to $\pi\in\mathcal{P}(Z\times Z)$ weakly; otherwise take a convergent subsequence. Estimating $\operatorname{EW}_{\lambda_{n}}$ using the plans $\tilde{\pi}_{n}\coloneqq(\operatorname{Id},\operatorname{Id})_{\sharp}\zeta_{n}$, where $\zeta_{n}\in\mathcal{P}(Z)$ is an arbitrary fixed-support barycenter, i.e., a minimizer of (11), we obtain
$\operatorname{EW}_{\lambda_{n}}^{2}(\mathbb{X}_{1},\mathbb{X}_{2})\leq\lambda_{n}\sum_{i=1}^{2}\operatorname{GW}^{2}(\mathbb{X}_{i},(Z,d_{Z},P_{i,\sharp}\tilde{\pi}_{n})=\lambda_{n}\operatorname{GWB}(\zeta).$ (50)
Thus, $\lambda_{n}\to 0$ implies $\operatorname{EW}_{\lambda_{n}}(\mathbb{X}_{1},\mathbb{X}_{2})\to 0$ and therefore $\operatorname{W}^{2}(P_{1,\sharp}\pi_{n},P_{2,\sharp}\pi_{n})\to 0$. Due to the stability of the Wasserstein distance (Ambrosio et al., 2005, Prop. 7.1.3.), the limit of $\pi_{n}$ has the form $\pi=(\operatorname{Id},\operatorname{Id})_{\sharp}\zeta$ for some $\zeta\in\mathcal{P}(Z)$. Moreover, we have
$\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},P_{1,\sharp}\pi_{n}))+\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},P_{2,\sharp}\pi_{n}))\leq\frac{1}{\lambda_{n}}\operatorname{EW}_{\lambda_{n}}^{2}(\mathbb{X}_{1},\mathbb{X}_{2})\leq\operatorname{GWB}(\mathbb{X}_{1},\mathbb{X}_{2}).$ (51)
Taking the limit $n\to\infty$, and exploiting (Beier & Beinert, 2025, Lem. I.1), we notice that the left-hand side becomes $\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},\zeta))+\operatorname{GW}^{2}(\mathbb{X}_{1},(Z,d_{Z},\zeta))$. Hence $\zeta$ has to be a fixed-support GW barycenter. $\Box$
Proof of Proposition 3.7
For arbitrary $\alpha\in\mathcal{P}(X_{1}\times Z_{1}\times Z_{2}\times X_{2})$, and for $\pi\coloneqq P_{Z_{1}\times Z_{2},\sharp}\alpha$ and $\gamma_{i}\coloneqq P_{X_{i}\times Z_{i},\sharp}\alpha$, the functional in (17) can be estimated by
$F_{\lambda}(\alpha)$ =${\iint_{(X_{1}\times Z_{1}\times Z_{2}\times X_{2})^{2}}^{}}\Bigl{[}\tfrac{1}{2}\bigl{(}d_{Z}^{2}(z_{1},z_{2})+d_{Z}^{2}(z_{1}^{\prime},z_{2}^{\prime})\bigr{)}+\lambda\sum_{i=1}^{2}(d_{X_{i}}(x_{i},x_{i}^{\prime})-d_{Z}(z_{i},z_{i}^{\prime}))^{2}\Bigr{]}\,\mathrm{d}\alpha(x_{1},z_{1},x_{2},x_{2})\,\mathrm{d}\alpha(x_{1}^{\prime},z_{1}^{\prime},z_{2}^{\prime},x_{2}^{\prime})$
$=\int_{Z_{1}\times Z_{2}}d_{Z}^{2}(z_{1},z_{2})\,\mathrm{d}\pi(z_{1},z_{2})+\lambda\sum_{i=1}^{2}\iint_{(X_{i}\times Z_{i})^{2}}(d_{X_{i}}(x_{i},x_{i}^{\prime})-d_{Z}(z_{i},z_{i}^{\prime}))^{2}\,\mathrm{d}\gamma_{i}(x_{i},z_{i})\,\mathrm{d}\gamma_{i}(x_{i}^{\prime},z_{i}^{\prime})$ (52)
$=\int_{Z_{1}\times Z_{2}}d_{Z}^{2}(z_{1},z_{2})\,\mathrm{d}\pi(z_{1},z_{2})+\lambda\left(G_{X_{1},Z}(\gamma_{1})+G_{X_{2},Z}(\gamma_{2})\right)\geq\operatorname{EW}^{2}_{\lambda}(\mathbb{X}_{1},\mathbb{X}_{2}).$ (53)
Now let $\pi^{*}$ and $\gamma_{i}^{*}$ be solutions of (13). Due to the gluing lemma (Villani, 2003, Lem. 7.6), there exists $\alpha^{*}\in\mathcal{P}(X_{1}\times Z_{1}\times Z_{2}\times X_{2})$ such that $P_{Z_{1}\times Z_{2},\sharp}\alpha^{*}=\pi^{*}$ and $P_{X_{i}\times Z_{i},\sharp}\alpha^{*}=\gamma_{i}^{*}$. For this $\alpha^{*}$, the inequality in (53) becomes sharp, i.e., $F_{\lambda}(\alpha^{*})=\operatorname{EW}_{\lambda}(\mathbb{X}_{1},\mathbb{X}_{2})$. Thus $\alpha^{*}$ solves (17). The other way round, let $\alpha^{*}$ be a solution of (17). As argued in the first part of the proof, the inequality in (53) has to be again sharp. Hence $\pi^{*}\coloneqq P_{Z_{1}\times Z_{2},\sharp}\alpha^{*}$ and $\gamma_{i}^{*}\coloneqq P_{X_{i}\times Z_{i},\sharp}\alpha^{*}$ solve (13). $\Box$